OpenStudy (motherhen544):

Write the equation of the line that passes through (−3, 5) and (2, 10) in slope-intercept form.

563blackghost (563blackghost):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\) \(\Large{By:thomaster}\) Welcome to Openstudy! And I hope you will enjoy this helpful site! ^^ Please check out the code of conduct to learn the rules of this site xD http://openstudy.com/code-of-conduct

563blackghost (563blackghost):

Heyo:) First we need to find the slope....so lets use slope formula.... \(\huge{\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\rightarrow \frac{10-5}{2-(-3)}}\)

OpenStudy (ajspeller):

Remember that y=mx+b slope is calculates as the difference in the y values divided by the x values The substitute in the values for the variables.

OpenStudy (motherhen544):

So that would make it 5/5 which is 1

563blackghost (563blackghost):

Correct :) Now we need to find the y-intercept....so we input the first point into the equation to find b.... \(\huge{y=x+b} \rightarrow 5=-3+b\) What would b equal?

OpenStudy (motherhen544):

2

563blackghost (563blackghost):

Not quite. We need to isolate b so we would add 3 to both sides since it is -3...what we do on one side we do to the other... \(\huge{5+3=(-3+3)b}\)

OpenStudy (motherhen544):

Sorry about that i was thinking something else I subtracted 3 from 5 instead of adding it. so b=7

563blackghost (563blackghost):

`b=8`. So now we know of the slope and y-intercept....In slope-intercept form the slope is m and the y-intercept is 8....so it would be set up as.... \(\huge{y=x+8}\)

OpenStudy (motherhen544):

Thank you so much sorry I can't add right now But you did help out a lot.

563blackghost (563blackghost):

Np :)